OpenStudy (anonymous):

240/100 in lowest terms and decimal

OpenStudy (anonymous):

240/100 = 24/10 = 12/5 = 2.4

OpenStudy (anonymous):

12/5

OpenStudy (anonymous):

2.4

OpenStudy (anonymous):

agree ^^

OpenStudy (anonymous):